Circular No. 8069

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003aj, 2003ak, 2003al, 2003am, AND 2003an Five additional apparent supernovae have been discovered, as tabu-

lated below. SNe 2003aj, 2003ak, and 2003al were reported by H. C. Ferguson, Space Telescope Science Institute, on behalf of the GOODS Treasury Team (+ K. S. Lee) and the Hubble Higher-z Supernova Team (cf. IAUC 8038), from images of the Chandra Deep Field South, each object detected by subtracting images taken on 2002 July 31–Aug. 4, Oct. 31–Nov. 3, and Dec. 19–22 from ACS images taken on Feb. 1–5, with each new object detected independently in each of four separate z-band (F850LP) exposures, and their z-band magnitudes tabulated below. SNe 2003am and 2003an were LOTOSS (cf. IAUC 8068) discoveries, reported by D. Weisz and W. Li from unfiltered KAIT images.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003aj Feb. 3.19 3^{h}32^{m}44.^{s}33 −27^{o}55′06′′.4 25.7 0′′.03 W, 0′′.08 S
2003ak Feb. 3.19 3 32 46.90 −27 54 49.3 25.8 0′′.28 E, 0′′.39 S
2003al Feb. 5.74 3 32 05.39 −27 44 29.2 24.5 0′′.03 W, 0′′.07 N
2003am Feb. 9.5 13 20 43.57 −22 02 53.5 17.8 2′′.9 W, 11′′.1 N
2003an Feb. 9.6 13 27 53.55 +28 30 29.2 16.9 4′′.2 E, 0′′.1 N

SN 2003al was detected also in two separate i-band (F775W) exposures, while SNe 2003aj and 2003ak were only marginally detected in the i band; none of the three HST objects was detected in images taken on 2002 Sept. 20.99, Nov. 2.51, and Dec. 22.10 UT (limiting magnitude z = 28). Additional approximate magnitude estimates: SN 2003am in ESO 576-G40, 2002 Jan. 12.3, [19.0; 2003 Feb. 10.4, 17.7. SN 2003an in MCG +05-32-22, Jan. 18.5, [18.5; Feb. 10.4, 17.0. SN 1997br also occurred in ESO 576-G40.

POSSIBLE NOVA IN M81 Weisz and Li also report a possible nova found on unfiltered KAIT

images from Feb. 9.4 (mag ∼ 17.8) and 10.4 UT (mag ∼ 17.9), with a hint of the new object on Feb. 6.4. The new object is located at α = 9^{h}55^{m}48^{s}.59, δ = +69^{o}03′04′′.3 (equinox 2000.0), which is 82′′.6 east and 50′′.8 south of the nucleus of M81 (= NGC 3031). A KAIT image taken on Jan. 29.4 showed nothing at this position (limiting mag ∼ 19.5). Weisz and Li add: “At the distance of M81 (3.6 Mpc; Freedman et al. 1994, Ap.J. 427, 628), the object has an absolute magnitude of ∼ −10, which is extremely bright for a nova. Spectroscopic observations are thus encouraged for the object.”

2003 February 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green